A geometric sequence can be used to describe the population of rabbits on a farm. The first spring the farmer purchased 10 rabbits. Four years later there are 810 rabbits at the farm. Assuming that none of the rabbits leave the​ farm, how many rabbits were on the farm in year 4​?
1
Expert's answer
2020-04-05T14:02:25-0400

First term in geometric progression is a= 10.

After 4 years a5= 810

The nth term is tn=a*rn-1.

So 10*r5-1=810,

r4=81,

r4=34.

Thus, r= 4.

Required is t4=10*34-1

= 10*33

=10*27

=270.

So there were 270 rabbits on the farm in the year 4.